OpenStudy (anonymous):

Kim rode her bike at a rate of 24km/h to the post office and 16km/h back home. If her trip to the post office and back took a total of 75 minutes, how far does Kim live from the post office?

OpenStudy (anonymous):

t = d/r Let d be the distance to the post office. Solve the following for d:\[\frac{ d}{24}+\frac{d}{16}=\frac{75}{60}\]d = 12 km